# Geometrical Aspects Of Resources Distribution In Quantum Random Circuits

### v1

@@ Abstract
Quantum random circuits , Non-locality , Quantum entanglement , CHSH-Mermin-Svetlichny inequalities
@@ 1 Introduction
Here we study the resources produced in different architectures of quantum computers using quantum random circuits (QRC) [8]. We focus on non-locality by using the Mermin [23] and the Svetlichny inequalities [29, 27]. We also study entanglement by appealing to different measures, such as tangle and concurrence (see B of this work). We first perform a theoretical study based on numerical simulations using Qiskit [26] and Amazon Braket SDK [4]. We analyze the performance of the devices with different levels of noise and number of shots, and compare universal (Clifford $+$ T) vs non-universal (Clifford) sets of elementary gates. Given a set of physical features which are thought of as a resource for quantum computing (as for example, non-locality), we quantify to which degree a device can generate a state space rich enough to generate the desired resources.
@@ 2 Quantum Random Circuits: distribution of resources
QRC are used in different areas of quantum information science [8]. In order to give a general idea, consider a scheme of circuit generation as depicted in Figure 1. On each stage of the protocol, random unitary gates are chosen and applied to the circuit. As a result, a random circuit is obtained. Remarkably, QRC are a promising tool to show quantum advantage [5]. Here, we use them to show how quantum resources are distributed among the possible states that a quantum computer is able to generate. An ideal quantum computer should be able to reach all possible states of a system of $N$-qubits. A real quantum device is able to implement a given set of elementary quantum gates on a fixed set of qubits. By generating QRC with different depths, sets of elementary gates, and levels of noise, we study how robust is the resources distribution with regard to the ideal case. Specifically, by comparing Clifford (H, S and CNOT) vs non Clifford (Clifford $+$ T) sets of gates, we highlight the differences of how quantum resources are distributed between universal vs non-universal quantum computing. In this section we focus on non-locality, as is expressed by the violation of certain inequalities. In the following section, we will deal with entanglement.
@@
where $M^{\prime}_{N-1}$ is obtained from $M_{N-1}$ by exchanging all the zero-indexed and and one-indexed $A$’s. The local realism limit is given by [3] (see also A):
@@
Consider the correlation functions $E(A_{1}^{(i)}A_{1}^{(j)}A_{3}^{(k)})$, which represent the expected value of the product of the measurement outcomes of single particle random variables $A_{1}^{(i)}$, $A_{2}^{(j)}$ and $A_{3}^{(k)}$ ($i,j,k$ are primed or not primed variables). Define
@@
It is important to remark that, in this work, the states generated using the Clifford set have a very different geometrical structure with regard to those generated with Clifford $+$ T. While in the latter the states generated tend to cover the entire set of quantum states (and it is very unlikely to obtain a given state twice), in the former, it is indeed likely to obtain repeated states, and they seem to be located in very restricted regions of the quantum state space. This explains, in part, the highly picked histograms displayed in what follows. Thus, for the Clifford set, the term "random" only applies to the way in which the gates of the elementary set are chosen to form each circuit (i.e., following the procedure indicated in Figure 1), and should not be confused with that of uniform distribution which applies when using random unitaries (and which the universal set resembles). The states thus generated with the Clifford set display a rich and intriguing geometrical structure that we will study in a separate work. Here, it suffices to have in mind that, in the histograms below, we have not deleted repeated states when computing the frequencies or counts, but deleting them doesn’t changes the essence of what we aim to show.
@@
The plots of Figure 4 clearly indicate how the resources distribution depend on the noise level. The comparison between universal vs non-universal sets of gates indicates that the main difference comes from the way in which resources are distributed. In both cases we obtain the maximum possible value of violation, which is attained by the GHZ state. This means that even a quantum device which is classically simulable can reach maximal non-locality. But, in a universal device, the different degrees of non-locality are homogeneously distributed among all possible values available in the full quantum state space. This is in line with previous works (compare with the results of [16] and [24]). By augmenting noise, one sees that the ideal histograms are deteriorated in both cases [^5] Notice that when noise is increased the histograms tend to a Gaussian-like distribution supported by non-violating values.. By comparing how an empirical histogram departs from the ideal case, one could certify how robust is the resources distribution in a concrete quantum device. The above analysis can be considered then as the basis of a certification protocol.
It is interesting to look at the violation fraction [^6] The violation fraction refers to the percentage of the states generated which violate a given inequality for some set of angles. as the number of qubits is increased (see Figure 5). For the Svetlichny inequality, we find that the violation fraction tends to zero as the number of qubits grows. This indicates that the relative volume of the set of genuine multipartite non-local states becomes very small as $N$ grows. In other words, the probability that a randomly picked pure state displays genuine non-local multipartite correlations, becomes very small in that limit. Contrarily, for the Mermin inequality, the violation percentage tends to remain very high as $N$ grows, indicating that the volume of the set of states displaying some form of non-local correlation is big (around $90\%$ of the whole state space). This is in agreement with the fact that the volume of the set of separable states tends to zero as $N$ goes to infinity. Notice that, given that entanglement does not necessarily entail non-locality, the results presented here are slightly different from those related to entanglement. In particular, our results indicate that the volume of the set of states that display no non-local correlation (at least with regard to the observables appearing in Mermin’s inequality) does not seem to tend to zero as $N$ grows. Here we are offering numerical evidence, but our findings give place to the question of how to prove theoretically that this is so. We will address all these questions in future works.
@@ 3 Entanglement measures and other quantifiers
Up to now, we have focused in non-locality, quantified as the violation levels of Mermin and Svetlichny inequalities. For the sake of comparison, it is also interesting to analyze what happens with entanglement. A short overview of the entanglement measures used in this work can be found in B. In Figure 9 we show how entanglement is distributed for different different entanglement measures and sets of elementary gates for three-qubits QRC.
@@
Figure 10: The various metrics of quantum correlations of the generated states are graphically represented along with the violation levels of the Svetlichny inequalities for three-qubit systems across 50,000 randomly generated circuits. The red dashed lines featured in the graphs correspond to the family of states studied in references [2, 6] (included here to facilitate a comparative analysis).
Figure 12: Distribution of Quantum Magic across 100,000 randomly generated circuits for: (A) three-qubit, (B) four-qubit and (C) five-qubit systems, featuring random depths. The circuits were constructed utilizing gates from the Clifford $+$ T set (left) and random unitaries, distributed in accordance with the Haar measure (right). Notably, for circuits constituted exclusively of the Clifford group, the quantum magic value is quantified as zero, hence its omission from the display. It is important to stress here the relevance of the depth of the circuits generated. One could fix a reasonable depth level chosen by inspection. Also, one can set the depth as a random parameter when generating the circuits, as we have done in the above plots. The later option yields results which resemble those of random unitaries with more precision.
@@ 4 Conclusions
In this work we have characterized the differences between the Clifford (Classically simulable) and Clifford $+$ T gates with regard to the violation of the Svetlichny and Mermin inequalities. Interestingly, even with a non-universal gate set, the maximal possible value of violation is obtained. Also, in the presence of noise, the Clifford set is still able to produce near maximal violations, as expected. The remarkable fact is that the main difference between the universal and the non-universal set of gates is that the latter fails to cover all possible values of the resource available in the quantum state space. This was clearly shown in Figure 8, with the presence of “holes´´ in the list of possible values of the resource. One could say that, essentially, the difference between universal vs non-universal quantum computing resides in the fact that, geometrically, the resources are distributed very differently along the quantum state space.
@@
Appendix A Explicit formulas for Mermin and Svetlichny inequalities
We show below explicit formulas for the Mermin polynomials from three to five qubit systems, $M_{2}$, $M_{3}$, $M_{4}$ and $M_{5}$:
$M_{2}=\frac{1}{2}\left(E\left(A_{1}A_{2}\right)+E\left(A_{1}A_{2}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}\right)\right),$
$M_{3}=\frac{1}{2}\left(E\left(A_{1}A_{2}A_{3}^{\prime}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}\right)\right),$
$M_{4}$ $=\frac{1}{4}(-E\left(A_{1}A_{2}A_{3}A_{4}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}\right)$
$+E\left(A_{1}A_{2}A_{3}A_{4}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}^{\prime}\right)$
$+E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}^{\prime}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}^{\prime}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}\right)$
$-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}\right)),$
$M_{5}$ $=\frac{1}{4}(-E\left(A_{1}A_{2}A_{3}A_{4}A_{5}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}A_{5}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}A_{5}\right)$
$+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}A_{5}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}A_{5}^{\prime}\right)$
$+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)+E\left(A_{1}A_{2}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)$
$-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)).$
We show below explicit formulas for the Svetlichny polynomials from four to five qubit systems, $S_{4}$ and $S_{5}$:
$S_{4}$ $=E\left(A_{1}A_{2}A_{3}A_{4}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}\right)-E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}\right)$
$-E\left(A_{1}A_{2}A_{3}A_{4}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}^{\prime}\right)$
$-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}^{\prime}\right)-E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}\right)$
$+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}^{\prime}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}\right),$
with a classical bound of $\left\langle S_{4}\right\rangle^{LR}\leq 8$ and a quantum bound of $\left\langle S_{4}\right\rangle^{QM}\leq 8\sqrt{2}$.
In the scenario involving five-qubits, the Svetlichny polynomial is formulated as follows:
$S_{5}$ $=E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}\right)+E\left(A_{1}A_{2},A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}A_{5}^{\prime}\right)$
$+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}A_{5}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)$
$+E\left(A_{1}A_{2}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)-E\left(A_{1}A_{2}A_{3}A_{4}A_{5}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)$
$-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}A_{5}\right)$
$-E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right)+E\left(A_{1}A_{2}A_{3}A_{4}A_{5}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}A_{5}^{\prime}\right)$
$-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}A_{4}A_{5}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}A_{5}\right)$
$+E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}A_{5}\right)+E\left(A_{1}A_{2}A_{3}A_{4}^{\prime}A_{5}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}A_{4}A_{5}^{\prime}\right)-E\left(A_{1}A_{2}A_{3}^{\prime}A_{4}^{\prime}A_{5}^{\prime}\right)$
$-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}A_{4}^{\prime}A_{5}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}A_{4}A_{5}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}A_{4}^{\prime}A_{5}^{\prime}\right),$
with a classical bound of $\left\langle S_{5}\right\rangle^{LR}\leq 16$ and a quantum bound of $\left\langle S_{5}\right\rangle^{QM}\leq 16\sqrt{2}$.
Appendix B A brief summary of the measures used in this work

### v2

@@ Abstract
Quantum random circuits $\cdot$ Quantum non-locality $\cdot$ Quantum entanglement $\cdot$ CHSH-Mermin-Svetlichny inequalities
@@ 1 Introduction
Here we study the resources produced in different architectures of quantum computers using quantum random circuits (QRC) [13]. We focus on non-locality by using the Mermin [14] and the Svetlichny inequalities [15, 16]. We also study entanglement by appealing to different measures, such as tangle and concurrence (see A of this work). We first perform a theoretical study based on numerical simulations using Qiskit [17] and Amazon Braket SDK [18]. We analyze the performance of the devices with different levels of noise and number of shots, and compare universal (Clifford $+$ T) vs non-universal (Clifford) sets of elementary gates. Given a set of physical features which are thought of as a resource for quantum computing (as for example, non-locality), we quantify to which degree a device can generate a state space rich enough to generate the desired resources.
@@ 2 Quantum Random Circuits: distribution of resources
QRC are used in different areas of quantum information science [13]. In order to give a general idea, consider a scheme of circuit generation as depicted in Figure 1. On each stage of the protocol, random unitary gates are chosen and applied to the circuit. As a result, a random circuit is obtained. Remarkably, QRC are a promising tool to show quantum advantage [1]. Here, we use them to show how quantum resources are distributed among the possible states that a quantum computer is able to generate. An ideal quantum computer should be able to reach all possible states of a system of $N$-qubits. A real quantum device is able to implement a given set of elementary quantum gates on a fixed set of qubits. By generating QRC with different depths, sets of elementary gates, and levels of noise, we study how robust is the resources distribution with regard to the ideal case.
Specifically, by comparing Clifford (H, S and CNOT) vs non Clifford (Clifford $+$ T) sets of gates, we highlight the differences of how quantum resources are distributed between universal vs non-universal quantum computing.
@@
In this section we focus on non-locality, as is expressed by the violation of certain inequalities. In the following section, we will deal with entanglement.
@@
where $M^{\prime}_{N-1}$ is obtained from $M_{N-1}$ by exchanging all the zero-indexed and and one-indexed $A$’s. The local realism limit is given by [20]:
@@
Consider the correlation functions $E(A_{1}^{(i)}A_{2}^{(j)}A_{3}^{(k)})$, which represent the expected value of the product of the measurement outcomes of single particle random variables $A_{1}^{(i)}$, $A_{2}^{(j)}$ and $A_{3}^{(k)}$ ($i,j,k$ are primed or not primed variables). Define
@@
It is important to remark that, in this work, the states generated using the Clifford set have a very different geometrical structure with regard to those generated with Clifford $+$ T. While in the latter the states generated tend to cover the entire set of quantum states (and it is very unlikely to obtain a given state twice), in the former, it is indeed likely to obtain repeated states (when the circuits are chosen at random). Thus, for the Clifford set, the term “random” only applies to the way in which the gates of the elementary set are chosen to form each circuit (i.e., following the procedure indicated in Figure 1), and should not be confused with that of uniform distribution which applies when using random unitaries (and which the universal set resembles). All possible states generated using Clifford gates can be actually computed, and the sets of states having the different possible values of degree of violation have the same weight as in the histograms displayed here. The states associated to the Clifford set display a rich and intriguing geometrical structure that we will study in a separate work.
@@
The plots of Figure 4 clearly indicate how the resources distribution depend on the noise level. The comparison between universal vs non-universal sets of gates indicates that the main difference comes from the way in which resources are distributed. In both cases we obtain the maximum possible value of violation, which is attained by the GHZ state. This means that even a quantum device which is classically simulable can reach maximal non-locality. But, in a universal device, the different degrees of non-locality are homogeneously distributed among all possible values available in the full quantum state space. This is in line with previous works (compare with [24]). By augmenting noise, one sees that the ideal histograms are deteriorated in both cases [^5] Notice that when noise is increased the histograms tend to a Gaussian-like distribution supported by non-violating values.. By comparing how an empirical histogram departs from the ideal case, one could certify how robust is the resources distribution in a concrete quantum device. The above analysis can be considered then as the basis of a certification protocol.
It is interesting to look at the violation fraction [^6] The violation fraction refers to the percentage of the states generated which violate a given inequality for some set of angles. as the number of qubits is increased (see Figure 5). For the Svetlichny inequality, we find that the violation fraction tends to zero as the number of qubits grows. This indicates that the relative volume of the set of genuine multipartite non-local states becomes very small as $N$ grows. In other words, the probability that a randomly picked pure state displays genuine non-local multipartite correlations, becomes very small in that limit. Contrarily, for the Mermin inequality, the violation percentage tends to remain very high as $N$ grows, indicating that the volume of the set of states displaying some form of non-local correlation is big (around $90\%$ of the whole state space). This is in agreement with the fact that the volume of the set of separable states tends to zero as $N$ goes to infinity. Notice that, given that entanglement does not necessarily entail non-locality, the results presented here are slightly different from those related to entanglement.
@@
In particular, our results indicate that the volume of the set of states that display no non-local correlation (at least with regard to the observables appearing in Mermin’s inequality) does not seem to tend to zero as $N$ grows. Here we are offering numerical evidence, but our findings give place to the question of how to prove theoretically that this is so. We will address all these questions in future works.
@@ 3 Entanglement measures and other quantifiers
Up to now, we have focused in non-locality, quantified as the violation levels of Mermin and Svetlichny inequalities. For the sake of comparison, it is also interesting to analyze what happens with entanglement. A short overview of the entanglement measures used in this work can be found in A. In Figure 9 we show how entanglement is distributed for different different entanglement measures and sets of elementary gates for three-qubits QRC.
@@
Figure 11: Distribution of Quantum Magic across 100,000 randomly generated circuits for three, four, and five-qubit systems, featuring random depths. The circuits were constructed utilizing gates from the Clifford $+$ T set and random unitaries, distributed in accordance with the Haar measure. Notably, for circuits constituted exclusively of the Clifford group, the quantum magic value is quantified as zero, hence its omission from the display. It is important to stress here the relevance of the depth of the circuits generated. One could fix a reasonable depth level chosen by inspection. Also, one can set the depth as a random parameter when generating the circuits, as we have done in the above plots. The latter option yields results which resemble those of random unitaries with more precision.
Figure 12: The various metrics of quantum correlations of the generated states are graphically represented along with the violation levels of the Svetlichny inequalities for three-qubit systems across 50,000 randomly generated circuits. The red dashed lines featured in the graphs correspond to the family of states studied in references [27] and [28] (included here to facilitate a comparative analysis).
4 Certifying quantum processors
The results obtained in the previous sections indicate that there is a sharp and quantifiable difference between, first, universal vs non-universal sets of gates, and second, ideal (noiseless) and non-ideal machines. By comparing the ideal vs experimentally obtained histograms of violation degrees, one can have a global measure of to what degree a quantum processor is able to generate the whole quantum state space. This suggests a methodology for testing the performance of a device, following the steps below:
  1. 1.
    Input: Specify the number $N$ of quantum random circuits generated, number of qubits to be used $Q$, depth of the circuit $D$ and number of shots $S$ to be used.
  2. 2.
    Implementation on the Device Using the native gates of the device, generate $N$ quantum random circuits ($S$ shots per circuit) and perform the measurements needed for the chosen resource to be tested (for the case, of a Bell-type inequality, we should perform the measurements related to the operator associated to that observable).
  3. 3.
    Comparison with ideal device: Compare the obtained results with the ideal ones (using the same set of gates). It could be to perform a histogram, compute the probabilities associated, and use some entropic measure to compare.
Implementation on quantum processors
Here we show an example of implementation in IonQ: Aria-1 and IQM: Garnet processors. Since implementing thousands of states in a real quantum processor would be expensive, here we focus on a variant of the steps mentioned in Methods and restrict ourselves to three-qubit systems. In the third step, instead of implementing all generated states, we proceed as follows. First, we divide the interval $[4,5.6]$ (the violation values range) in subintervals of length $0.2$. Next, from each interval $I$, we look at the set $\Delta_{I}$ of generated states that have a maximal violation value in the interval $I$. From the set $\Delta_{I}$, we pick a representative state minimizing the number of gates and the number of entangling gates (CZ in the IQM: Garnet and MS in the IonQ: Aria-1 case).
Figure 13: Schematic representation of IQM Garnet’s connectivity and the spin orientations measured in one of our experiments. Each arrow represents a possible spin direction of qubits $1$, $2$, and $4$. All possibilities are listed in the right.
In order to generate the random circuits, we modified the function used in the previous section to generate circuits of variable length, and connectivity of each processor and using only native gates (PRX and CZ for IQM: Garnet, and GPi, GPi2 and MS for IonQ: Aria-1). Local spin observables are implemented as proper rotations with PRX or GPi-GPi2 gates depending on the processor (the theoretical angles obtained for the synthetic data can be easily translated in terms of those gates). A schematic picture of the IQM: Garnet connectivity and the different spin directions used in a test of a Svetlichny violation is shown in Figure 13. The results of running ten states are displayed in Figure 15, together with their corresponding theoretical values of maximal violations.
It is also instructive to see what happens when the same violation experiment is repeated several times. In Figure 15 we show a plot with ten repetitions for each state of Figure 15. The results obtained indicate that while a violation of the Svetlichny inequality is clearly obtained, the results are far from the ideal case (compare these results with those of Figure 3).
Figure 14: Comparison between IonQ: Aria-1 (green dots) and IQM: Garnet (red dots). For each interval, we chose a representative state which uses a minimal number of MS and CZ gates respectively (and a minimum global number of gates). For each of them, we store the values of angles of spin measurement of (theoretically) maximal violation, in which one instance of experiment is chosen for each QPU.
Figure 15: Degree of Svetlichny violation for different instances of three qubits random circuits implemented in IQM Garnet. We have prepared ten repetitions of each state to study the fluctuation in the different runs of the experiment.
As the results of each instance of a Svetlichny violation fluctuate, we have performed an analysis of the dependence with the number of shots. The results are displayed in Figure 16. Clearly, the fluctuation tends to decrease as the number of shots grow. But using $1000$ shots gives a good idea of the degree of violation attained by the QPU in the mean.
Before we go on, it is important to make a digression about the interpretation of the results presented in Figure 15. When trying to generate a state $\rho$ in a quantum computer, one needs to implement a series of instructions $I=\{i_{1},i_{2},....,i_{n}\}$ in the quantum processing unit.
Figure 16: We have prepared ten repetitions of each state.
Let us call the state resulting from those instructions $\rho^{ideal}_{I}$. If there were no imperfections present during the process, then, we should have $\rho=\rho^{ideal}_{I}$. But in an actual experiment, $\rho$ and $\rho^{ideal}_{I}$ might be significantly different. Therefore, the implementation of the instructions is automated in such a way to optimize the task of creating a state as similar as possible to $\rho$. For that reason, the actual instructions implemented in the QPU will be denoted by a list $I^{\prime}=\{i_{1}^{\prime},i_{2}^{\prime},....,i_{n^{\prime}}^{\prime}\}$. While $I^{\prime}$ might be quite different from $I$, if the optimization process is successful, the resulting state $\rho^{real}_{I^{\prime}}$ will be very similar to $\rho$. In other words, we end up with $\rho^{real}_{I^{\prime}}\approx\rho^{ideal}_{I}$. Notice that the notation $\rho^{real}_{I^{\prime}}$ refers to the actual state generated in the QPU (which might be in turn different from the ideal state $\rho^{ideal}_{I^{\prime}}$ produced with noiseless instructions $I^{\prime}$). The situation is similar when we implement measurements. Suppose that in order to emulate a certain measurement in a QPU we need to implement the list of instructions $M=\{m_{1},m_{2},...,m_{l}\}$. Let us call the resulting ideal state $\rho^{ideal}_{I+M}$ (i.e., the state that results from implementing noiseless instructions). Again, in an actual experiment, we would obtain $\rho^{real}_{I^{\prime}+M^{\prime}}$. Ideally, to perform the desired measurement in the state $\rho$, we should run the circuit with the instructions $I\cup M$ and sample in the computational basis. As performing a measurement in a basis different than the computational one is equivalent to performing a suitably chosen rotation and measuring in the computational basis, if the optimization process works reasonably well, one concludes that $\rho^{real}_{I^{\prime}+M^{\prime}}$ is the best representative of the ideal process of preparing $\rho$ and measuring $M$. Notice also that in two different runs of the instructions $I\cup M$, one might obtain different circuits and, in principle, the chosen qubits might be even different. One could say that this is not relevant, because the goal is to reproduce the different copies of $\rho^{real}_{I+M}$ with the highest possible fidelity.
Figure 17: Degree of violation for GHZ states in Aria-1 and Garnet for two to five qubits. The statistics for the correlations of each shot are computed with 1000 shots.
It is important to notice that the above discussion can be applied mutatis mutandis to every quantum experiment, since every quantum experiment actually performed can be reduced to a list of operations on a physical system. The quantum systems used might be different (for example, different instances of photons or atoms), and the implementation of the procedures and measurements might slightly differ from one repetition to another. But the main goal of every experiment is to generate several copies of the ideal state and measurements (which are as similar as possible to it), in order to perform the required statistics.
Using IonQ’s, IQM’s transpilation system and IQM: FakeAdonis simulator we performed a final experiment by sending GHZ states. The results are presented in Figure 17.
5 Genuine multipartite non-locality vs contextuality
The experimental test of Bell inequalities was challenging due to the presence of possible loopholes: defects in the experimental implementation can be used to give explanations of the observed data in which the local-realism hypothesis can be maintained. After several years of huge efforts, the so-called loophole-free Bell tests were finally announced [29, 30, 31]. Taking into account how hard it was to reach that experimental degree of certification, one may wonder up to which extent it is realistic to affirm that genuine multipartite non-locality was produced in a given quantum processor. Indeed, in NISQ devices, qubits are usually very close, and the measurement of the different parties seem to be far from the conditions required for a loophole-free non-locality test. A quick thought indicates that it is unlikely that anyone can construct in the near future a quantum processor in which genuine multipartite non-locality can be tested in a loophole-free way. Therefore, we should be conservative about the certification of non-locality in current quantum processors (see also the discussion in [32], page 2). Still, the violation of Svetlichny inequalities in quantum processors can be considered as a form of non-classicality, given that it discards a certain form of hidden variable models. Perhaps, it is safer to say that they represent a particular form of contextuality in the following sense. For a certain family of observables (the local spin orientations with suitably chosen angles), there exists no global probability distribution satisfying certain constrains. Alternatively, one possible strategy is to work under the no-cross-talk assumption [32]. In any case, it is reasonable to consider the violations presented in the previous section as a form of non-classicality, given that they imply that the state presents genuine multipartite entanglement.
The above observations give place to a very interesting debate about the foundations of quantum physics that has technological implications. A proponent of the thesis that loophole-free certified non-locality is a necessary and sufficient condition for quantum advantage, will have to face the problem that it is extremely unlikely that a quantum processor will be ever able to give it for granted. This has implications on the development strategy of QPUs since, in the quest for quantum advantage, it would be necessary to move forward under the assumption that non-locality is actually present. Otherwise, efforts should be devoted to testing this feature in a loophole-free way, but this doesn’t seem realistic with the extant technologies and experimental techniques. On the contrary, a defender of the thesis that quantum advantage is all about contextuality will still have a point in moving forward. One could always consider a quantum processor as a simulator of genuine multipartite non-locality that uses contextuality as a resource. Is this enough for quantum advantage? Only time can tell the answer to this intriguing question. We will address this fascinating subject with more detail in a future work. With this short section, we just want to draw the reader’s attention on the relevance of this problem, since it connects both, foundations of quantum physics debates, and technological developments. We think that the results presented in this work contribute to a better understanding of the challenges that the development of quantum processors must face.
@@ 6 Conclusions
In this work we have characterized the differences between the Clifford (Classically simulable) and Clifford $+$ T gates with regard to the violation of the Svetlichny and Mermin inequalities. Interestingly, even with a non-universal gate set, the maximal possible value of violation is obtained. Also, in the presence of noise, the Clifford set is still able to produce near maximal violations, as expected. The remarkable fact is that the main difference between the universal and the non-universal set of gates is that the latter fails to cover all possible values of the resource available in the quantum state space. This was clearly shown in Figure 8, with the presence of “holes” in the list of possible values of the resource. One could say that, essentially, the difference between universal vs non-universal quantum computing resides in the fact that, geometrically, the resources are distributed very differently along the quantum state space.
@@
CODE AVAILABILITY
The code used to simulate the QRC is available at https://github.com/QuantumUnica/QRC_Search
Appendix A A brief summary of the measures used in this work